OpenStudy (anonymous):

The circulation of a newsletter decreased from 3,200 to 2,464. What was the percent of decrease in circulation? A. 23% B. 77% C. 2.3% D. 129%

OpenStudy (anonymous):

If you write it as rate equivalences you got something like: 100/3200=x/2464 solving x you can find: x=(100∗2464)/3200 x=77 We've found that only 77% of newsletters are in circulation, the percent of decrease is: 100−77=23

OpenStudy (anonymous):

Wait so I did this problem but I got 77% and that was wrong. but I guess I just forgot that last step! so the answer is 23 right?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Alternatively, subtract beginning value 3200 from end value 2464 to obtain the decrease (a negative quantity), and divide that by 3200, and multiply the result by 100%.

OpenStudy (anonymous):

woa math why is your ss thing a square

OpenStudy (anonymous):

Gees thanks you guys are a life saver!!

OpenStudy (mathmale):

a square? Please explain where that came from.

OpenStudy (anonymous):

your smartscore . the 65 is in a rhombus

OpenStudy (mathmale):

Oh, that?!?! You're first to notice. I get the privilege of seeing ad-free Open Study screens because I've signed up to contribute $10/month. OpenStudy and you guys are worth it. :)

OpenStudy (anonymous):

oh . nice :)